I need 2 gallons of paint @ $22.95 each, with 8% sales tax...What is the total?
I have already done the rest of this riduculously long problem but my brain is fried!
Thanks so much!

Answer by jim_thompson5910(28476)

One gallon is $22.95, so two gallons is 22.95 + 22.95 = 45.90

A shortcut is 2*22.95 = 45.90


So ignoring tax, the total cost is $45.90
-------------------------------------------------------
Now let's add on the 8% tax.

We do this by first taking 8% of the total 45.90 to get

8% of 45.90 = 0.08*45.90 = 3.672 = 3.67 (note: I'm rounding to the nearest penny)

So the amount in taxes you pay is $3.67

Now add this result to the total:

45.90 + 3.67 = 49.57

So in total (with taxes now), the final total is $49.57